# Taxonomy of amorphous ternary phase diagrams: the importance of interaction parameters

### v1

@@ Abstract
Phase diagrams offer a comprehensive representation of the thermodynamics of multi-component systems. However, for ternary amorphous systems, the different possible phase diagram types and their existence conditions are still unclear. Based on the systematic screening of the parameter space for three amorphous ternary material systems representative of polymer, small molecule, and solvent materials, we report identifying twenty-one phase diagram types generated using Flory-Huggins theory. The proposed classification relies on the number of immiscible material pairs, distinct miscibility gaps, and three-phase regions. We infer existence rules by mapping the type of phase diagram to the range of interaction parameters in the three-dimensional parameter space. Depending on the number of immiscible pairs (0, 1, 2, or 3), four well-known phase-diagram types are found to be very likely. However, numerous uncommon phase diagrams are observed within a small parameter window around the critical interaction parameter values. Regarding the processability window, the size of the miscible region becomes sensitive to interaction parameter variations, mainly when they are close to critical values. The sensitivity decreases for materials with increasing molar size. Finally, the paper includes a successful comparison of simulated phase diagrams with experimental data, showcasing the real-world relevance of this theoretical analysis and its applicability for optimizing solution processing methods efficiently.

### v2

@@ Abstract
Understanding phase diagrams is essential for material selection and design, as they provide a comprehensive representation of the thermodynamics of mixtures. This work delivers a broad and systematic overview of possible ternary phase diagrams for amorphous systems representative of polymers, small organic molecules, and solvents. Thanks to computationally efficient methods, an unprecedented library of $>$80,000 ternary phase diagrams is generated based on a systematic screening of interaction parameters. Twenty-one phase diagram types, including unreported ones, are identified. They are classified according to simple rules related to the number of immiscible material pairs, of miscibility gaps, and of three-phase regions. They are mapped onto the three-dimensional interaction parameters space, providing a clear picture of their likelihood and existence conditions. Four well-known phase-diagram types with 0, 1, 2, or 3 immiscible pairs are found to be the most likely. The numerous uncommon phase diagrams are mostly observed within a small parameter window around the critical interaction parameter values. For the most common phase diagram types, we show that the size of the processability window becomes sensitive to interaction parameter variations close to critical values. The sensitivity decreases for materials with increasing molar size. Finally, successful comparisons of simulated and experimental phase diagrams showcase the real-world relevance of this theoretical analysis. The presented results lay a robust foundation for rational design of solution processing conditions and for blend morphology control. Immediate applications include organic thin films and the identification of green solvents for sustainable processing.